OpenStudy (luigi0210):

Integration:

OpenStudy (luigi0210):

\[\LARGE \int~\frac{cosx}{\sqrt{4-cos^2x}}~dx\]

OpenStudy (anonymous):

can you help me? i @ted you

OpenStudy (anonymous):

\[\cos^2x = 1-\sin^2x\] \[\int\limits \frac{ cosx }{ \sqrt{\sin^2x+3} }dx \] u = sinx, du = cosxdx \[\int\limits \frac{ 1 }{ \sqrt{u^2+3} }du\] \[u=\sqrt{3}tant~~~~du = \sqrt{3}\sec^2t dt\] Finish it off

OpenStudy (luigi0210):

Thank you batman :D

OpenStudy (anonymous):

Yw

OpenStudy (anonymous):

can you guys help me now?

geerky42 (geerky42):

Once again Openstudy is saved!

OpenStudy (anonymous):

@geerky42 can you help me